### v2

# The Structure of the Majorana Clifford Group

Abstract
In quantum information science, Clifford operators and stabilizer codes play a central role for systems of qubits (or qudits). In this paper, we study the analogous objects for systems of Majorana fermions. A crucial role is played by fermion parity symmetry, which is an unbreakable symmetry present in any system in which the fundamental degrees of freedom are fermionic. We prove that the subgroup of parity-preserving fermionic Cliffords can be represented by the orthogonal group over the binary field $\mathbb{F}_{2}$, and we show how it can be generated by braiding operators and used to construct any (even-parity) Majorana stabilizer code. We also analyze the frame potential for this so-called p-Clifford group, proving that it is equivalent to the frame potential of the ordinary Clifford group acting on a fixed-parity sector of the Hilbert space.
I Introduction
The standard model of quantum computation is formulated in terms of commuting qubit degrees of freedom, but one can also formulate a model of quantum computation in terms of anti-commuting fermionic degrees of freedom [1]. In the standard model, an important role is played by the Clifford group, which is a special subgroup of the full unitary group whose elements act in a simple way on Pauli operators [2]. Here we study the fermionic analog of this structure, the Majorana Clifford group, which is a special subgroup of unitaries whose elements act in a simple way on Majorana fermion operators. And while any Majorana Clifford turns out be just a Jordan-Wigner transformation away from becoming a Pauli Clifford, the extra ingredient of fermion superselection places new constraints on which operations can be considered to be physical.
One motivation for an analysis of the fermionic Clifford group is provided by a promising class of fault-tolerant quantum computer architectures that employs the topological braiding of Majorana fermion modes to encode logical information [3, 4, 5, 6, 7, 8]. While their topological structure provides these systems with some inherent degree of protection against outside noise, the concatenation of additional error-correcting (stabilizer) codes generated by fermionic Clifford operations acting on these Majorana modes will likely still be necessary to achieve adequate macroscopic resilience. Fermionic stabilizer codes, which were initially proposed by Bravyi et al. [9], have since garnered a rich literature [10, 11, 12, 13, 14, 15]. Such codes have advantages compared to their Pauli-based counterparts, most notably their natural protection against parity-switching errors thanks to fermionic superselection rules.
Another motivation arises from the Sachdev-Ye-Kitaev (SYK) model which has recently become one of the most-studied physical systems in the fields of condensed matter and quantum gravity [16, 17, 18, 19, 20]. Describing a large number of randomly coupled Majorana fermions, it manages to provide a simple toy model for holography, connecting concepts in quantum information and high energy physics. Recently, we proposed a tensor network model of the SYK model and analyzed a qubit-based toy model employing random Clifford gates [21]. The desire to generalize this toy model to fermionic degrees of freedom was the starting point of this work, although we have deferred any analysis of the corresponding fermionic tensor networks to future work. Random sampling from the fermionic Clifford group could also be of use for questions concerning fermionic random circuits and fermionic classical shadows [22].
Given these motivations and the broadly important role played by the ordinary Clifford group in quantum information and physics, this paper presents a thorough analysis of the fermionic Clifford group. We are primarily concerned with the description and manipulation of systems composed of $2n$ Majorana fermion modes $\chi_{i}$, $i=1,\ldots,2n$ obeying
$\chi_{i}^{\dagger}=\chi_{i},\quad\{\chi_{i},\chi_{j}\}=2\,\delta_{ij},$ (1)
and equipped with a fundamental representation on a $2^{n}$-dimensional Hilbert space. Fermion superselection is encoded via their parity symmetry, which is generated by the Hermitian operator
$(-1)^{F}\propto\prod_{i=1}^{2n}\chi_{i}.$ (2)
Fermion parity is important because any physical fermionic observable or circuit must commute with it. Hence, we focus on p-Cliffords, the subgroup of Majorana Cliffords which preserve fermion parity. Our findings show that the group of p-Cliffords admits a projective representation as binary orthogonal matrices, a subgroup of the symplectic group describing all Cliffords. Each orthogonal matrix can be generated by a small set of Householder reflections, which in the Hilbert space representation corresponds to the p-Cliffords being generated by Majorana braiding operators. Using a canonical basis for even-parity stabilizers, we show that by, adding ancilla modes, any other even stabilizer basis (and its corresponding logical operators) can be generated by applying an appropriate p-Clifford. Last but not least, we prove that the p-Clifford group can be considered to be a unitary design on a fixed-parity sector of the Hilbert space, but like the ordinary Clifford group only manages to be a 3-design.
Overall, the rest of this paper is structured as follows: In Section II we formally define Majorana strings, prove some basic properties, and discuss the the connections between operator weight, fermion parity and the resulting superselection rules. We also highlight the relation to Pauli strings and the resulting Hilbert (sub-)space representation. Next, in Section III we formally define the general (fermionic) Clifford group, and identify the p-Cliffords as its subgroup, characterized in terms of orthogonal matrices over the field $\mathbb{F}_{2}$ and efficiently generated by braiding operators. In Section IV we discuss fermionic stabilizer states and how, with the help of an ancilla systems, they can be generated using p-Cliffords. Fourth, in Section V we discuss the problem of sampling from the p-Clifford group, analyze the frame potential, and show that the p-Clifford is a 3-design within sectors of fixed fermion parity. We conclude with an outlook and discussion of open problems in Section VI.
Note: At the time of writing, the authors became aware of independent research being conducted on the same topic [15]. Their results inspired some of the terminology in Section III.3, and a variation of their stabilizer generation algorithm using our established framework in section IV.4.
II Majorana Strings
To describe strings of fundamental Majorana fermion operators and the superoperators that map between them, it is most convenient to use the language of symplectic projective representations over the binary field $\mathbb{F}_{2}$. This construction follows analogous (and equivalent) programs applied to ordinary Pauli strings [23].
II.1 General Properties
A reasonable definition for Hermitian strings made of some even number of elementary Majorana mode operators is as follows:
Definition 1 (Hermitian Majorana String).
Given $2n$ independent Majorana fermion operators $\chi_{i}$ satisfying (1), we define a general Hermitian Majorana string $\operatorname{\mu}(v)$ as a normal-ordered product of said operators, such that
$\operatorname{\mu}(v)\coloneqq(i)^{v^{T}\operatorname{\operatorname{\omega}_{L}}v}\cdot\chi_{1}^{v_{1}}\chi_{2}^{v_{2}}\cdots\chi_{2n}^{v_{2n}},\,v\in\mathbb{F}_{2}^{2n},$ (3)
where we implicitly defined the lower-triangular matrix
$\omega_{L}\coloneqq\begin{pmatrix}0&0&\cdots&0&0\\ 1&0&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 1&1&\cdots&0&0\\ 1&1&\cdots&1&0\end{pmatrix}\in\mathbb{F}_{2}^{2n\times 2n}.$ (4)
For the sake of completeness we also identify its upper-triangular counterpart as $\operatorname{\operatorname{\omega}_{U}}\equiv\operatorname{\operatorname{\omega}_{L}}^{T}$.
It is straightforward to show that this definition indeed satisfies $\operatorname{\mu}(v)^{\dagger}=\operatorname{\mu}(v)$ by reordering the operators after taking the adjoint. All expressions involving elements of the finite field $\mathbb{F}_{2}\equiv(\{0,1\},+,\cdot)$ or vector spaces over $\mathbb{F}_{2}$ are understood to be evaluated mod 2.
With this we can define the equivalent of the Pauli group as the span of all (not necessarily Hermitian) Majorana strings:
Definition 2 (Majorana Group).
The Majorana group $\operatorname{\mathcal{M}}_{2n}$ is the set of all Majorana strings of length $2n$:
$\operatorname{\mathcal{M}}_{2n}\coloneqq\{(i)^{a}\cdot\operatorname{\mu}(v)\,|\,a\in\mathbb{Z}_{4},v\in\mathbb{F}_{2}^{2n}\}.$ (5)
$\operatorname{\mathcal{M}}_{2n}$ is a non-abelian group under operator multiplication, and an irreducible projective representation of the abelian group $\mathbb{Z}_{4}\times\mathbb{F}_{2}^{2n}$ (see Lemma 2).
The group structure of $\operatorname{\mathcal{M}}_{2n}$ follows directly from the composition property of two (Hermitian) Majorana strings
$\operatorname{\mu}(v)\operatorname{\mu}(v^{\prime})=\zeta(v,v^{\prime})\cdot\operatorname{\mu}(v+v^{\prime}).$ (6)
Here $\zeta(v,v^{\prime})$ is some non-trivial [^1] The reason for these cumbersome expressions is the incompatibility between $\mathbb{F}_{2}$ having addition mod 2 and $(i)^{a}=(i)^{a+4}$ implying addition mod 4. The same issue also plagues qubit Pauli strings. coefficient that can be shown to be of the form
$\zeta(v,v^{\prime})\equiv(-1)^{v^{T}\operatorname{\operatorname{\omega}_{L}}v^{\prime}+f(v,v^{\prime})}\,(i)^{v^{T}\omega v^{\prime}},$ (7)
where we also identified the symmetric function $f(v,v^{\prime})=f(v^{\prime},v)$:
$\begin{split}f(v,v^{\prime})\equiv{}&(v^{T}\omega_{L}v)(v^{\prime T}\omega_{L}v^{\prime})\\ &+v^{T}\omega v^{\prime}(v^{T}\omega_{L}v+v^{\prime T}\omega_{L}v^{\prime}+1).\end{split}$ (8)
Both expressions $\zeta(v,v^{\prime})$ and $f(v,v^{\prime})$ involve a new symmetric [^2] Analogous matrices over fields of higher order are actually skew-symmetric, but for $\mathbb{F}_{2}$ this makes no difference. matrix $\operatorname{\omega}$ which arises in the derivation as the sum of $\operatorname{\operatorname{\omega}_{L}}$ and its transpose $\operatorname{\operatorname{\omega}_{U}}$:
$\operatorname{\omega}\coloneqq\operatorname{\omega}_{L}+\operatorname{\omega}_{U}=\begin{pmatrix}0&1&\cdots&1&1\\ 1&0&\cdots&1&1\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 1&1&\cdots&0&1\\ 1&1&\cdots&1&0\end{pmatrix}\equiv I^{c}.$ (9)
Here $I^{c}$ denotes the matrix complement of the identity.
Definition 3 (Complement).
For any $n\times m$ matrix $A$ over $\mathbb{F}_{2}$, we define its matrix complement $A^{c}$ as the matrix one gets when exchanging $0\leftrightarrow 1$. This is equivalent to writing
$A^{c}\equiv J-A,$ (10)
where $J$ is the $n\times m$ matrix with all its entries equal to 1. Vector complements are defined analogously using the corresponding all-ones vector:
$j\coloneqq(1,1,\ldots,1)^{T}.$ (11)
While (7) and (8) are highly nontrivial expressions, it is straightforward to argue (and explicitly prove) that
$\zeta(v,v)=1,$ (12)
since $\mu(v)$ is Hermitian with ±$\pm 1$ eigenvalues, and therefore satisfies $\mu(v)^{2}=I\equiv\mu(0)$.
II.2 (Anti-) Commutation
Despite the complicated composition law (6) for two Majorana strings, things simplify considerably if we consider just their commutation:
$\operatorname{\mu}(v)\operatorname{\mu}(v^{\prime})=(-1)^{v^{T}\operatorname{\omega}v^{\prime}}\operatorname{\mu}(v^{\prime})\operatorname{\mu}(v)$ (13)
The expression $v^{T}\operatorname{\omega}v^{\prime}$ therefore indicates if two strings commute or anti-commute. And since any operator commutes with itself, it also has to satisfy
$v^{T}\operatorname{\omega}v=0\quad\forall v\in\mathbb{F}_{2}^{2n}.$ (14)
It’s straightforward to show that this is indeed the case, e.g. using (20) in Section II.3. With this property of $\operatorname{\omega}$ we can hence make the following statement:
Lemma 1 (Symplectic Product).
The map $\left\langle\cdot,\cdot\right\rangle:\mathbb{F}_{2}^{2n}\times\mathbb{F}_{2}^{2n}\rightarrow\mathbb{F}_{2}$ acting according to
$\left\langle v,v^{\prime}\right\rangle\coloneqq v^{T}\operatorname{\omega}v^{\prime}$ (15)
defines a symplectic product, i.e. it is bilinear, alternating ($\left\langle v,v\right\rangle=0$) and non-degenerate. $\operatorname{\omega}$ is therefore called a symplectic form.
Proof.
[…]
$\begin{split}\lambda_{1}&=N-1\mod{2},\quad e_{1}=(1,1,\ldots,1)^{T}\\ \lambda_{i}&=1,\quad e_{i}=(0,\ldots,0,\underbrace{1,1}_{\mathclap{\text{$(i-1)$th and $i$th position}}},0,\ldots,0)^{T},\quad 1<i\leq N.\end{split}$ (16)
But since we chose $N=2n$ to be even, we have $\lambda_{1}=1$ and the spectrum is therefore not degenerate. ∎
II.3 Weight & Parity
As explained in Section I, two important properties of a Majorana string $\operatorname{\mu}(v)$ (and the associated vector $v$) are its (operator) weight and (fermion) parity.
Definition 4 (Operator Weight & Fermion Parity).
  • The (operator) weight $\operatorname{w}(\operatorname{\mu}(v))\equiv\operatorname{w}(v)$ of a Majorana string $\operatorname{\mu}(v)$ is defined as the number of (elementary) non-identity Majorana operators $\chi_{i}$ it is composed of. Equivalently, the weight is equal to the number of non-zero entries in the string’s vector representation $v\in\mathbb{F}_{2}^{2n}$.
  • The (fermion) parity $\operatorname{p}(\operatorname{\mu}(v))\equiv\operatorname{p}(v)$ of a Majorana string $\operatorname{\mu}(v)$ can be defined as the string’s operator weight mod 2. Equivalently, the parity can be defined in terms of the dot product of the string’s vector representation with itself or $j$ (11) (over $\mathbb{F}_{2}$):
    $\operatorname{p}(v)\equiv v^{T}v=v^{T}j.$ (17)
    If $\operatorname{p}(v)=1$, $\operatorname{\mu}(v)$ (or $v$) is said to have odd parity. If $\operatorname{p}(v)=0$, $\operatorname{\mu}(v)$ (or $v$) is said to have even parity.
The concept of operator weights is of general interest in the context of stabilizer codes used for quantum error correction. This is because the weight indicates the minimum number of elementary operations one would have to perform in an experimental setup to apply a single Majorana string to a given system. Applying strings with large weights thus requires precise control of the system and a significant amount of time. This allows for errors to occur during the probing, which can therefore be hard to detect. In general, one therefore desires to minimize the weights of the strings being applied as much as possible.
The parity of Majorana strings is of physical relevance due to the constraints imposed on it by fermion superselection. In general, superselection rules dictate that for an observable $O$ (including the Hamiltonian) to be physical, it must obey
$\langle\psi_{1}|O|\psi_{2}\rangle=0$ (18)
for any two states $\lvert\psi_{1}\rangle$, $\lvert\psi_{2}\rangle$ from different superselection sectors. This implies that states from different sectors can’t be put into coherent superpositions, and that the superselection-defining observable is conserved.
In the context of (Majorana) fermions, the two relevant sectors correspond to even and odd parity respectively, and the conserved observable is hence the parity operator
$(-1)^{F}\equiv\mu(j),$ (19)
which is uniquely defined up to an overall sign. Superselection then dictates that any (Hermitian) Majorana string describing a physical observable must commute with $(-1)^{F}$. However, using the general property
$\operatorname{\omega}v=\operatorname{p}(v)\cdot j-v=\begin{cases}v,&\operatorname{p}(v)=0\\ v^{c},&\operatorname{p}(v)=1\end{cases},$ (20)
one comes to the conclusion that $v^{T}\operatorname{\omega}j=v^{T}j=\operatorname{p}(v)$ for all $v\in\mathbb{F}_{2}^{2n}$ per definition (17), and thus the the commutation property (13) implies
$(-1)^{F}\operatorname{\mu}(v)\,(-1)^{F}=(-1)^{\operatorname{p}(v)}\cdot\operatorname{\mu}(v).$ (21)
Therefore only the strings with even parity commute and hence can act as physical observables or elements of a stabilizer. This does however not imply that odd-parity strings are without use, as they can for example describe logical operators acting on stabilizer states. This is discussed in more detail in Section IV.3.
Besides their physical relevance, even Majorana strings also have the useful property that their parity is preserved under composition with other even strings. Hence it makes sense to identify the subspace of all even-parity vectors in $\mathbb{F}_{2}^{2n}$:
$\mathbb{F}_{2}^{2n,+}\coloneqq\{v\in\mathbb{F}_{2}^{2n}|\operatorname{p}(v)=0\},$ (22)
and the associated even-parity Majorana subgroup:
$\operatorname{\mathcal{M}}^{+}_{2n}\coloneqq\{(i)^{a}\cdot\operatorname{\mu}(v)\,|\,a\in\mathbb{Z}_{4},v\in\mathbb{F}_{2}^{2n,+}\}\leq\operatorname{\mathcal{M}}_{2n}.$ (23)
One can easily see that the same can’t be done for Majorana strings with odd parity due to it not being invariant under composition. Nevertheless, it is useful to define their respective counterparts as the sets $\mathbb{F}_{2}^{2n,-}$ and $\operatorname{\mathcal{M}}_{2n}^{-}$.
II.4 Connection to Pauli Strings
Let’s connect what we have established so far to the analogous representation of (qubit) Pauli strings. In that context a binary vector of length $2n$ can be mapped to a tensor product of $n$ elementary Pauli operators [23]:
$\operatorname{\pi}(v)=(i)^{v^{T}\operatorname{\operatorname{\eta}_{L}}v}\,Z^{v_{1}}X^{v_{2}}\otimes\cdots\otimes Z^{v_{2n-1}}X^{v_{2n}},$ (24)
where $Z$ and $X$ are the standard Pauli matrices and $\operatorname{\operatorname{\eta}_{L}}$ is the equivalent of $\operatorname{\operatorname{\omega}_{L}}$, here corresponding to the lower-triangular part of the matrix
$\operatorname{\eta}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}^{\oplus\,n}.$ (25)
Analogous to $\operatorname{\omega}$, $\eta$ encodes the commutation structure of Pauli strings and hence defines a symplectic product. We can therefore find an invertible linear transformation $W\in\operatorname{GL}(2n,\mathbb{F}_{2})$ that maps between these two symplectic forms:
$\omega=W^{T}\,\eta\,W.$ (26)
This has the implication that if two Majorana strings $\operatorname{\mu}(v),\operatorname{\mu}(v^{\prime})$ (anti-)commute, then the corresponding Pauli strings $\pi(Wv),\pi(Wv^{\prime})$ also do:
$\left\langle v,v^{\prime}\right\rangle_{\operatorname{\mu}}=\left\langle Wv,Wv^{\prime}\right\rangle_{\operatorname{\pi}}$ (27)
What $W$ therefore describes is the $\mathbb{F}_{2}$-representation of a Jordan-Wigner transformation between operator product strings of Majorana fermions, and tensor product strings of Pauli operators:
$\operatorname{\mu}(v)\cong\operatorname{\pi}(Wv).$ (28)
This is trivially true for $n=1$ since in that case $\operatorname{\omega}=\operatorname{\eta}$, which coincides with the natural identification $Z\cong\chi_{1}$, $X\cong\chi_{2}$. In general, $W$ can be shown to be [^3] Plugging (30) into the left side of (26) gives $(\operatorname{\eta}^{c})_{L}\operatorname{\eta}\,(\operatorname{\eta}^{c})_{U}=\operatorname{\eta}\,(\operatorname{\eta}^{c})_{L}(\operatorname{\eta}^{c})_{U}=\operatorname{\eta}\operatorname{\eta}^{c}=\operatorname{\eta}\operatorname{\eta}\operatorname{\omega}=\operatorname{\omega},$ (29) where we used that $(\operatorname{\eta}^{c})_{L}$ commutes with $\operatorname{\eta}$, $\operatorname{\eta}^{c}=(\operatorname{\eta}^{c})_{L}(\operatorname{\eta}^{c})_{U}$, $\operatorname{\eta}^{c}=\operatorname{\eta}\operatorname{\omega}$ because $\operatorname{\eta}$ has odd-parity rows and columns, and $\operatorname{\eta}^{2}=I$.
$W=(\operatorname{\eta}^{c})_{U}=\begin{pmatrix}1&0&1&1&\cdots&1&1\\ 0&1&1&1&\cdots&1&1\\ 0&0&1&0&\cdots&1&1\\ 0&0&0&1&\cdots&1&1\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&0&\cdots&1&0\\ 0&0&0&0&\cdots&0&1\\ \end{pmatrix},$ (30)
which is the upper-triangular part of the matrix complement of $\operatorname{\eta}$ (including the diagonal). It is also easy to see that $W$ is its own inverse:
$W^{2}=I,$ (31)
meaning that it maps both vector representations of Majorana strings to ones of Pauli strings, and vice versa. From the structure of (30) it also becomes apparent that low-weight strings in one basis are generally mapped to high-weight strings in the other. This is to be expected, as the most basic Jordan-Wigner transformations are used to assign local fermionic degrees of freedom to topological domain walls in 1D Ising models.
II.5 Hilbert Space Representation
So far we have not mentioned a concrete representation of Majorana strings on a Hilbert space $\operatorname{\mathcal{H}}$. Fortunately, thanks to the identification with Pauli strings we can simply choose the one defined implicitly in terms of (28) and (24). This means that the trace of a Majorana string (also referred to as the character of the chosen representation) is proportional to
$\operatorname{tr}(\operatorname{\mu}(v))=\operatorname{tr}(\pi(Wv))=2^{n}\delta_{v,0},$ (32)
i.e. all strings except for multiples of the identity are traceless, and the dimension of $\operatorname{\mathcal{H}}$ is $2^{n}$.
Lemma 2 (Irreducibility).
The Hilbert space representation of Majorana strings defined in terms of (28) and (24) is irreducible.
Proof.
It is a known result of representation theory (see e.g. [24]) that it suffices for the averaged sum over the absolute square of the characters $\operatorname{tr}((i)^{a}\operatorname{\mu}(v))$ to be equal to 1, which is indeed the case due to (32):
$\frac{1}{|\mathbb{Z}_{4}|}\frac{1}{|\mathbb{F}_{2}^{2n}|}\sum_{a\in\mathbb{Z}_{4}}\sum_{v\in F^{2n}}|\operatorname{tr}((i)^{a}\operatorname{\mu}(v))|^{2}=1.$ (33)
II.6 Majorana Algebras
While taking the trace of a Majorana string (or any operator that can be expanded in terms of Majorana strings) is straightforward, the lack of a tensor product structure in (3) means that the Hilbert space has no obvious factorization with regard to the individual mode operators $\chi_{i}$, and hence no way to take a partial trace. Instead one has to adopt the notion of Majorana (sub-) algebras [25], which we briefly review here due to their importance e.g. for analyzing the entanglement and entropy of fermionic subsystems.
Any operator $O$ acting on the Hilbert space $\operatorname{\mathcal{H}}$ that can be expressed as a linear combination of Majorana strings such that
$O\equiv\sum_{v\in\mathbb{F}_{2}^{2n}}c_{v}\cdot\operatorname{\mu}(v),\quad c_{v}\in\mathbb{C},$ (34)
is an element of the (fundamental) Majorana algebra. In fact, it is straightforward to show that a Majorana algebra composed of $2n$ modes encompasses all complex-valued $2^{n}\times 2^{n}$ matrices. Using (32) and the fact that Hermitean strings square to the identity, the coefficients $c_{v}$ of the expansion can be determined using:
$c_{v}=\frac{1}{2^{n}}\operatorname{tr}(O\operatorname{\mu}(v)).$ (35)
A Majorana subalgebra $\mathcal{A}$ is any algebra spanned by Majorana strings which only contain an (ordered) even subset $A\subseteq I$ of all modes $I=\{1,2,\ldots,2n\}$ (here represented by their respective indices). These strings are therefore of the form
$\operatorname{\mu}(\Pi(v_{A}\oplus 0_{\bar{A}})),\quad v_{A}\in\mathbb{F}_{2}^{2n_{A}},$ (36)
where $n_{A}=|A|/2$, and $\Pi$ is some fixed permutation matrix. Note however that we run into issues with regard to the Hilbert space representation of the subalgebra. The strings as depicted in (36) are still endowed with the ordinary $2^{n}\times 2^{n}$-representation (32), even though the algebra itself is isomorphic to the $2^{n_{A}}\times 2^{n_{A}}$ algebra of complex matrices. Hence this representation is reducible, which is especially problematic when considering properties like entropy that intrinsically depend on the Hilbert space dimension. To compensate, we introduce a new irreducible representation $\operatorname{\mu}_{A}(v_{A})$ which satisfies (3) and (32) with $n\rightarrow n_{A}$. The reduced operator $O_{A}$ corresponding to the chosen subalgebra is then given by
$\begin{split}O_{A}&=\sum_{v_{A}}c_{v_{A}}\cdot\operatorname{\mu}_{A}(v_{A}),\\ c_{v_{A}}&=\frac{1}{2^{n}}\operatorname{tr}(O\operatorname{\mu}(\Pi(v_{A}\oplus 0_{\bar{A}}))).\end{split}$ (37)
III The Fermionic Clifford Group
Having laid out the basic principles of Majorana strings, it is now time to consider the possible ways to transform (between) them. Quantum mechanics dictates that any such transformation should take the form of a unitary (super-) operator, and the largest class of such operators that map Majorana strings to other Majorana strings is called the (Majorana) Clifford group:
Definition 5 (Majorana Clifford Group).
The Majorana Clifford group $\operatorname{\mathcal{C}}_{2n}$ is the group of all unitary automorphisms acting on the Majorana group $\operatorname{\mathcal{M}}_{2n}$ (5), i.e. all unitary operators $\operatorname{\gamma}$ that map (general) Majorana strings to other such strings:
$\operatorname{\gamma}\operatorname{\mu}(v)\operatorname{\gamma}^{\dagger}\equiv c(v)\cdot\operatorname{\mu}(S(v))$ (38)
for some functions $c:\mathbb{F}_{2}^{2n}\rightarrow\operatorname{\mathbb{C}}$ and $S:\mathbb{F}_{2}^{2n}\rightarrow\mathbb{F}_{2}^{2n}$ that depend on $\operatorname{\gamma}$.
Understanding the structure of the Clifford group is therefore equivalent to understanding the allowed options for $S$ (and to lesser extent $c$).
III.1 Symplectic Structure
Unitarity generally preserves the (anti-) commutation relations of operators $O_{1},O_{2}$ in the sense that
$[UO_{1}U^{\dagger},UO_{2},U^{\dagger}]=U[O_{1},O_{2}]U^{\dagger}.$ (39)
The same must therefore be true for elements of the Clifford group. Since it was shown in section II.2 that two Majorana strings $\operatorname{\mu}(v),\operatorname{\mu}(v^{\prime})$ (anti-) commuting is equivalent to the symplectic product $\left\langle v,v^{\prime}\right\rangle$ being zero (or one), the function $S$ associated to a Clifford element as described in definition 5 has to preserve the structure of said product. We later prove that any $S$ that does so is necessarily linear and hence symplectic:
Definition 6 (Symplectic Group).
The (binary) symplectic group $\operatorname{Sp}(2n)\equiv\operatorname{Sp}(2n,\mathbb{F}_{2})$ is the group of all invertible linear transformations $S\in\operatorname{GL}(2n,\mathbb{F}_{2})$ that preserve the symplectic product:
$\left\langle Sv,Sv^{\prime}\right\rangle=\left\langle v,v^{\prime}\right\rangle\quad\forall\,v,v^{\prime}\in\mathbb{F}_{2}^{2n}.$ (40)
This definition does not depend on any specific choice of basis, meaning it applies equally to the Majorana and Pauli case. In the basis of former, (40) is equivalent to
$S^{T}\omega S=\omega\quad\forall S\in\operatorname{Sp}(2n).$ (41)
Obtaining the associated group element in the Pauli basis (where one would require $S^{T}\eta S=\eta$ instead) is then as simple as performing the Jordan-Wigner transformation (30):
$S\rightarrow WSW.$ (42)
Proposition 1 (Group Order of $\operatorname{Sp}(2n)$).
The total number of elements of the symplectic group $\operatorname{Sp}(2n)$ is [26]
$\lvert\operatorname{Sp}(2n)\rvert=2^{n^{2}}\prod_{i=1}^{n}(4^{i}-1).$ (43)
A method to uniformly and efficiently sample a symplectic matrix $S_{i}$ (in the Pauli basis) by associating it to a unique integer $1\leq i\leq\lvert\operatorname{Sp}(2n)\rvert$ is illustrated in [27].
With that we can formalize and further expand upon the previous statements to fully classify the elements of the (Majorana) Clifford group $\operatorname{\mathcal{C}}_{2n}$:
Theorem 1 (Symplectic Representation of Cliffords).
  1. 1.
    For $\operatorname{\gamma}$ as in (38) to be an element of the Majorana Clifford group $\operatorname{\mathcal{C}}_{2n}$, the associated mapping $S$ must be linear and therefore symplectic. Conversely, for any symplectic matrix $S$ there is at least one $\operatorname{\gamma}\in\operatorname{\mathcal{C}}_{2n}$ that acts accordingly.
  2. 2.
    If $\operatorname{\gamma}(S)$ is a Clifford with associated symplectic $S$, then (ignoring overall phases) $\operatorname{\mu}(a)\operatorname{\gamma}(S)$, $a\in\mathbb{F}_{2}^{2n}$ is also a Clifford with the same associated symplectic $S$. Conversely, any two Clifford group elements $\operatorname{\gamma}(S),\operatorname{\gamma}^{\prime}(S)$ are equivalent up to some Majorana string $\operatorname{\mu}(a)$ and a phase:
    $\operatorname{\gamma}(S)\sim\operatorname{\gamma}^{\prime}(S)=e^{i\phi}\operatorname{\mu}(a)\operatorname{\gamma}(S).$ (44)
  3. 3.
    Let $[\operatorname{\gamma}]:\operatorname{Sp}(2n)\rightarrow\operatorname{\mathcal{C}}_{2n}/\sim$ be the mapping that assigns to each symplectic $S$ the corresponding equivalence class $[\operatorname{\gamma}(S)]$ of Cliffords. Then this mapping defines a representation of $\operatorname{Sp}(2n)$:
    $[\operatorname{\gamma}(S)]\,[\operatorname{\gamma}(T)]=[\operatorname{\gamma}(ST)].$ (45)
A proof for this discrete variation of the Stone-von Neumann Theorem is found in Appendix A.1.
Note that it is not possible to find a closed form expression for the coefficients $c(v)$ in (38). This is due to the field $\mathbb{F}_{2}$, which describes Majorana (and qubit Pauli) strings, being special from a mathematical perspective. Still, by evaluating $\operatorname{\gamma}(S)\operatorname{\mu}(v)\operatorname{\mu}(v^{\prime})\operatorname{\gamma}^{\dagger}(S)$ in two different ways one can obtain the following relation:
$c(v)\,c(v^{\prime})=\frac{\zeta(v,v^{\prime})}{\zeta(Sv,Sv^{\prime})}\cdot c(v+v^{\prime}),$ (46)
which implies that $c(v)^{2}=c(0)=1$ due to (12), as expected from unitarity. Stronger statements can be made for qudit Pauli strings with $d>2$ prime [23], but since the overall phase of a quantum state has no physical meaning, we can safely ignore it most of the time, and focus solely on the symplectic representation $S$. $\operatorname{\gamma}(S)$ is therefore used to denote any Clifford that acts accordingly.
Corollary 1 (Order of the Clifford Group).
The size of the quotient set $\operatorname{\mathcal{C}}_{2n}/\operatorname{U}(1)$, i.e. the number of elements in $\operatorname{\mathcal{C}}_{2n}$ that (ignoring phases) are pairwise distinct, is
$\lvert\operatorname{\mathcal{C}}_{2n}/\operatorname{U}(1)\rvert=2^{n^{2}+2n}\prod_{i=1}^{n}(4^{i}-1).$ (47)
Proof.
Theorem 1 directly implies that the quotient set size factorizes according to
$|\operatorname{\mathcal{C}}_{2n}/\operatorname{U}(1)|=|\operatorname{Sp}(2n)|\cdot|\mathbb{F}_{2}^{2n}|.$ (48)
Plugging in (43) and $|\mathbb{F}_{2}^{2n}|=2^{2n}$ then results in what was to be shown. ∎
III.2 Parity-Preserving Cliffords
So far the Pauli and Majorana Clifford groups are mathematically equivalent. But unlike Pauli strings, Majorana strings have the additional extrinsic property of fermion parity. It is easy to see that Clifford operators do not obey parity superselection in general, e.g.
$\begin{pmatrix}1&0\\ 1&1\end{pmatrix}\begin{pmatrix}1\\ 0\end{pmatrix}=\begin{pmatrix}1\\ 1\end{pmatrix}$ (49)
depicts a symplectic matrix (Clifford operator) mapping between vectors (Majorana strings) of different parity. For an element $\operatorname{\gamma}(S)$ to respect the parity of any string, the associated symplectic $S$ must satisfy
$p(Sv)=(Sv)^{T}(Sv)=v^{T}(S^{T}S)v\stackrel{{\scriptstyle!}}{{=}}v^{T}v=p(v)$ (50)
for all $v\in\mathbb{F}_{2}^{2n}$, and therefore be orthogonal:
Definition 7 (Orthogonal Group).
The (binary) orthogonal group $\operatorname{O}(2n)\equiv\operatorname{O}(2n,\mathbb{F}_{2})$ is the group of all invertible linear transformations $A\in\operatorname{GL}(2n,\mathbb{F}_{2})$ that preserve the canonical inner product:
$A^{T}A=AA^{T}=I.$ (51)
In fact, any orthogonal matrix is also symplectic, meaning that $\operatorname{O}(2n)$ is a subgroup of the symplectic group $\operatorname{Sp}(2n)$ [^4] In the Majorana basis this is easy to see, as any $A\in\operatorname{O}(2n)$ can only have odd-parity row and column vectors, leading to $A^{T}\operatorname{\omega}A=A^{T}A^{c}=A^{T}A\operatorname{\omega}=\operatorname{\omega}.$ (52) :
$\operatorname{O}(2n)\leq\operatorname{Sp}(2n).$ (53)
This also implies that the order of $\operatorname{O}(2n)$ must divide the order of $\operatorname{Sp}(2n)$ (43), which is true since:
Proposition 2 (Group Order of $\operatorname{O}(2n)$).
The total number of elements of the orthogonal group $\operatorname{O}(2n)$ is [26]
$\begin{split}\lvert\operatorname{O}(2n)\rvert&=2^{n^{2}}\prod_{i=1}^{n-1}(4^{i}-1)\\ &=\lvert\operatorname{Sp}(2n)\rvert/(4^{n}-1).\end{split}$ (54)
Overall, we can therefore identify a subgroup of the Clifford group $\operatorname{\mathcal{C}}_{2n}$ that acts according to $\operatorname{O}(2n)$ and hence preserves parity:
Corollary 2.
(Parity-Preserving Clifford Subgroup) The Majorana Clifford group $\operatorname{\mathcal{C}}_{2n}$ admits a parity-preserving subgroup $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$. Its elements are of the form $\operatorname{\mu}(a)\operatorname{\gamma}(S)$ with $S\in\operatorname{O}(2n)$ and $a\in\mathbb{F}_{2}^{2n,+}$. The size (mod phases) of this so-called p-Clifford group is therefore
$\begin{split}|\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}/\operatorname{U}(1)|&=|\operatorname{O}(2n)|\cdot|\mathbb{F}_{2}^{2n,+}|\\ &=2^{n^{2}+2n-1}\prod_{i=1}^{n-1}(4^{i}-1).\end{split}$ (55)
Note that we have $\operatorname{\mathcal{M}}_{2n}^{+}\leq\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$, because while odd-parity strings don’t affect the parity of other strings when acting from both sides, they can indeed change the parity of states in the Hilbert space. This is shown explicitly in Section IV.2. Overall, all elements of $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ hence commute with $(-1)^{F}$ (19) because
$\operatorname{\gamma}(S)\operatorname{\mu}(j)\operatorname{\gamma}^{\dagger}(S)=c(j)\cdot\operatorname{\mu}(Sj)=\operatorname{\mu}(j),$ (56)
where $Sj=j$ due to all $S\in\operatorname{O}(2n)$ having odd-parity rows, and $c(j)=1$ using arguments made in Appendix A.2.
III.3 Generating Set
For the ordinary Pauli Clifford group, the Gottesman-Knill theorem [28] states (among other things) that $\operatorname{\mathcal{C}}_{2n}$ can be efficiently generated from repeated applications of tensor-products containing CNOT, Hadamard, and the phase gate (or an equivalent minimal set) as the only non-trivial operators. In the symplectic representation, this is equivalent to the fact that any element $S\in\operatorname{Sp}(2n)$ can be decomposed into a product of symplectic transvections
$\tau_{a}:v\mapsto v+\left\langle a,v\right\rangle_{\operatorname{\pi}}a,\quad a\in\mathbb{F}_{2}^{2n}.$ (57)
Additionally, all transvections can be written as products and (in the Pauli basis [^5] In the Majorana basis it is straightforward to show that the direct sum of symplectics is not necessarily symplectic anymore. More specifically, only direct sums of orthogonal matrices are again orthogonal and therefore symplectic.) direct sums of just 3 independent ones:
  • For $n=1$, there are two such independent transvections (e.g. for $s=(1,0)^{T}$ and $h=(1,1)^{T}$), from which all other symplectics can be generated by repeated multiplication of $\tau_{s}$ (which corresponds to the phase gate) and $\tau_{h}$ (which corresponds to the Hadamard).
  • For $n=2$, it can be shown that there is only one independent transvection $\tau_{c}$ (for e.g. $c=(1,0,0,1)^{T}$) since any other one can be written as $\tau_{c}$ times some direct sum of $n=1$ symplectics. In fact, CNOT can be shown to correspond to the symplectic $(\tau_{s}\oplus\tau_{h}\tau_{s}\tau_{h})\,\tau_{c}$ in the Pauli basis.
  • For $n>2$ no other independent transvections exist.
We now present an analogous statement for just the p-Clifford subgroup $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$. Our focus hereby is only on generating the orthogonal action since adding a phase change-inducing element of $\operatorname{\mathcal{M}}^{+}_{2n}$ is trivial.
Theorem 2 (Gottesman-Knill for p-Cliffords).
Any element of the p-Clifford group $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}\leq\operatorname{\mathcal{C}}_{2n}$ (mod $\operatorname{\mathcal{M}}^{+}_{2n}$, and excluding $\operatorname{BRAID}_{2n}(j)$ due to (56)) can be efficiently generated from products of $\operatorname{BRAID}_{2}$ and $\operatorname{BRAID}_{4}$ operators, which for any $k\geq 0$ are defined as:
$\begin{split}\operatorname{BRAID}_{k}(a)&\coloneqq\exp\left(i\frac{\pi}{4}\operatorname{\mu}(a)\right)\\ &=\frac{1}{\sqrt{2}}\big{(}I+i\operatorname{\mu}(a)\big{)},\end{split}$ (58)
with $a\in\mathbb{F}_{2}^{2n},\,\operatorname{w}(a)=k$ (compare to [13, 15]).
Proof.
We first note the well-known fact that the orthogonal group $\operatorname{O}(2n)$ is generated by the set of Householder reflections [26]
$h_{a}:v\mapsto v+(a^{T}v)\,a,\quad a\in\mathbb{F}_{2}^{2n,+}.$ (59)
Compare this to the action of $\operatorname{BRAID}_{k}(a)$ on a general Majorana string $\operatorname{\mu}(v)$, which can be shown to be
$\begin{split}&\operatorname{BRAID}_{k}(a)\operatorname{\mu}(v)\operatorname{BRAID}^{\dagger}_{k}(a)\\ ={}&\frac{1}{2}\left(1+(-1)^{\left\langle a,v\right\rangle}\right)\operatorname{\mu}(v)\\ &+\frac{i}{2}\left(1-(-1)^{\left\langle a,v\right\rangle}\right)\operatorname{\mu}(a)\operatorname{\mu}(v)\\ \equiv{}&(i\,\zeta(a,v))^{a^{T}v}\cdot\operatorname{\mu}(v+(a^{T}v)\,a),\end{split}$ (60)
using the fact that $\left\langle a,v\right\rangle=a^{T}v$ for $\operatorname{p}(a)=0$ and any $v$. The symplectic representations of the (even-parity) $\operatorname{BRAID}_{k}$ operators are hence the Householder reflections:
$\operatorname{BRAID}_{k}(a)\equiv\operatorname{\gamma}(h_{a}),$ (61)
which implies that they generate the orthogonal action of $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ [^6] This argument can be extended to the whole Clifford group by also including braiding operators with odd weights. In fact, symplectic transvections reduce to Householder reflections when restricting $a$ in (57) to be even.. In appendix B we provide an algorithm that can generate any orthogonal matrix from Householder reflections in $\mathcal{O}(n^{3})$ steps.
What is left to show is that all reflections (except $h_{j}$) can be generated from just the ones corresponding to strings with weight 2 and weight 4. This can be done in at most $\mathcal{O}(n^{3})$ steps by noting that one can write $h_{a}=I+aa^{T}$, and hence
$h_{b}h_{a}h_{b}=h_{h_{b}a}.$ (62)
Given some reflection associated to a vector $a$ with weight $k>4$ even, the problem therefore reduces to finding weight 4 reflections that map it to a vector with weight 2 or 4. This is always possible because any weight 4 Householder is of the form $\operatorname{\omega}_{4}\oplus I_{2n-4}$ (up to permutations), essentially allowing us to use (20) with regard to any 4 (odd-parity) entries of $a$ to reduce its weight by 2. This reduction has to be performed $\mathcal{O}(k)$ times, each step involving two $\mathcal{O}(n^{2})$ matrix applications due to (62). ∎
IV Fermionic Stabilizers
We now discuss how to define and construct general fermionic stabilizers in terms of Majorana operator strings, such that obey the parity superselection rules.
Definition 8 (Fermionic Stabilizer).
A $[[n,k]]_{f}$ (Majorana) fermionic stabilizer $\operatorname{\mathcal{S}}$ is an abelian subgroup of the even-parity Majorana group $\operatorname{\mathcal{M}}^{+}_{2n}$ that is generated by $r=n-k\leq n$ independent strings and does not contain the operator $-I$.
We don’t mention explicit examples of fermionic stabilizer codes as they can be readily found in e.g. [9, 15, 12, 10, 11].
IV.1 Symplectic Representation
The reason why a stabilizer can have at most $n$ independent generators becomes apparent when one considers the symplectic equivalent:
Definition 9 (Isotropic Subspaces).
A subspace $M$ of the vector space $\mathbb{F}_{2}^{2n}$ is said to be isotropic iff
$\left\langle m,m^{\prime}\right\rangle=0\quad\forall m,m^{\prime}\in M.$ (63)
The largest possible isotropic subspaces have dimension $n$, and are called maximally isotropic [23].
Due to fermion superselection we only care about isotropic subspaces of the even-parity subspace $\mathbb{F}_{2}^{2n,+}$. In this case each maximally isotropic subspace if of the form
$M_{\text{max}}\equiv SM_{\text{std}},\quad S\in\operatorname{O}(2n),$ (64)
where $M_{\text{std}}$ is the canonical maximally isotropic subspace given in terms of its basis as row vectors of a $n\times 2n$ matrix
$M_{\text{std}}\sim\begin{pmatrix}1&1&0&0&0&\cdots&0&0&0\\ 0&0&1&1&0&\cdots&0&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ 0&0&0&0&0&\cdots&0&1&1\end{pmatrix}.$ (65)
It is easy to check that one can’t add more rows that are orthogonal with regard to the already present ones.
Each element of an (even-parity) isotropic subspace $M\subset\mathbb{F}_{2}^{2n,+}$ of dimension $r=n-k$ can be associated to a Majorana string of (at least one) $[[n,k]]_{f}$ fermionic stabilizer $\operatorname{\mathcal{S}}$. Fixing the basis of $M$ to be $\{b_{1},\ldots,b_{k}\}$, one possible choice for this mapping is given by
$\operatorname{\mathcal{S}}\sim\operatorname{\bar{\operatorname{\mu}}}(M)\coloneqq\{\operatorname{\bar{\operatorname{\mu}}}(m)\,|\,m\in M\},$ (66)
where we defined
$\operatorname{\bar{\operatorname{\mu}}}(m)\coloneqq\prod_{i=1}^{k}\operatorname{\mu}(b_{i})^{m_{i}},\quad m=\sum_{i=1}^{k}m_{i}b_{i}\in M.$ (67)
Using the modified (Hermitian) Majorana string $\operatorname{\bar{\operatorname{\mu}}}(m)$ instead of $\operatorname{\mu}(m)$ is beneficial because the resulting effective phases factors of the stabilizer strings are compatible with composition:
$\operatorname{\bar{\operatorname{\mu}}}(m)\operatorname{\bar{\operatorname{\mu}}}(m^{\prime})=\operatorname{\bar{\operatorname{\mu}}}(m+m^{\prime})\quad\forall m,m^{\prime}\in M,$ (68)
which is only possible due to the strings $\operatorname{\mu}(b_{i})$ commuting pairwise. Any other possible stabilizer associated to $M$ differs from $\operatorname{\bar{\operatorname{\mu}}}(M)$ only by the signs in front of each element $\operatorname{\bar{\operatorname{\mu}}}(m)$ [23]. These signs naturally must also be compatible with composition (i.e. form a character). More precisely:
Theorem 3.
Any $[[n,k]]_{f}$ (Majorana) fermionic stabilizer $\operatorname{\mathcal{S}}$ is of the form
$\operatorname{\mathcal{S}}(M,v)=\{(-1)^{v^{T}m}\operatorname{\bar{\operatorname{\mu}}}(m)\,|\,m\in M\},$ (69)
where $M\subset\mathbb{F}_{2}^{2n,+}$ is an even-parity isotropic subspace of dimension $r=n-k$, and $v$ is an element of the quotient set $\mathbb{F}_{2}^{2n}/M$  [^7] A priori, we allow $v$ to have odd parity, as it only affects the leading phases. We will return to this in section IV.2..
IV.2 Stabilizer Spaces & States
Using the elements of a stabilizer $\operatorname{\mathcal{S}}(M,v)$, we can construct a projection operator onto the subspace of logical states, i.e. those states that have unit eigenvalue:
Lemma 3.
Given a $[[n,k]]_{f}$ (Majorana) fermionic stabilizer $\operatorname{\mathcal{S}}(M,v)$, then the space of states $\lvert\psi\rangle\in\operatorname{\mathcal{H}}$ satisfying
$(-1)^{v^{T}m}\operatorname{\bar{\operatorname{\mu}}}(m)\lvert\psi\rangle=\lvert\psi\rangle\quad\forall m\in M$ (70)
is $2^{k}$-dimensional, and projected onto by the operator
$\Pi(M,v)=\frac{1}{|M|}\sum_{m\in M}(-1)^{v^{T}m}\operatorname{\bar{\operatorname{\mu}}}(m).$ (71)
Proof.
It is straightforward to check that $\Pi(M,v)^{2}=\Pi(M,v)$ due to $M$ forming a group under addition. Hence the dimension of the space it projects onto is
$\operatorname{tr}(\Pi(M,v))=\frac{2^{n}}{|M|}=2^{n-r}=2^{k}.$ (72)
Finally, the projection operator satisfies
$(-1)^{v^{T}m}\operatorname{\bar{\operatorname{\mu}}}(m)\Pi(M,v)=\Pi(M,v)\quad\forall m\in M,$ (73)
and therefore only projects onto the logical states. ∎
In particular, for the case $r=n$ (i.e. $M$ being maximally isotropic), we can identify a single (logical) state $\lvert M,v\rangle$ satisfying (70), such that
$\Pi(M,v)\equiv\lvert M,v\rangle\langle M,v\rvert,\quad\dim(M)=n.$ (74)
The set $\{\lvert M,v\rangle\,|\,v\in\mathbb{F}_{2}^{2n}/M\}$ forms a basis of $\operatorname{\mathcal{H}}$.
Note that even though the projection operators are composed of even-parity stabilizer elements, they don’t necessarily project onto Hilbert space states that have even parity. Indeed, the eigenvalue of $(-1)^{F}\equiv\operatorname{\bar{\operatorname{\mu}}}(j)$ [^8] We can write $(-1)^{F}$ as $\operatorname{\bar{\operatorname{\mu}}}(j)$ because $j$ is guaranteed to be an element of $M$ if it is maximally isotropic. associated to a stabilizer state $\lvert M,v\rangle$ is
$\langle(-1)^{F}\rangle=\operatorname{tr}(\operatorname{\bar{\operatorname{\mu}}}(j)\,\lvert M,v\rangle\langle M,v\rvert)=(-1)^{\operatorname{p}(v)}.$ (75)
Applying a Majorana string $\operatorname{\mu}(a)$ to the state so that
$\operatorname{\mu}(a)\lvert M,v\rangle\langle M,v\rvert\operatorname{\mu}(a)=\lvert M,v+a\rangle\langle M,v+a\rvert$ (76)
therefore changes the parity of the state if $\operatorname{p}(a)=1$, even though the parity of the stabilizer elements is unaffected. This further illustrates why only the even-parity Majorana group $\operatorname{\mathcal{M}}^{+}_{2n}$ is a subgroup of $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$, and not $\operatorname{\mathcal{M}}_{2n}$.
IV.3 Logical Operators
When the fermionic stabilizer is not maximal, i.e. the space of logical states $\operatorname{\mathcal{H}}_{\text{logic}}$ has dimension $12^{k}>$ 1, one can identify Majorana string operators that map between those states while respecting the overall structure. This leads to the following definition:
Definition 10 (Logical Operators).
The logical operators $\operatorname{\mathcal{L}}$ associated to a fermionic stabilizer $\operatorname{\mathcal{S}}$ are the unitary elements of $\operatorname{\mathcal{M}}_{2n}\setminus\operatorname{\mathcal{S}}$ that commute with all stabilizer elements (but not necessarily among themselves).
One can convince oneself that $\operatorname{\mathcal{L}}$ acting on the logical Hilbert space $\operatorname{\mathcal{H}}_{\text{logic}}$ is generated by $2k$ independent Majorana strings, analogously to how there are $2n$ independent Majorana strings for the physical Hilbert space $\operatorname{\mathcal{H}}$. Furthermore, those logical operators can have odd parity iff ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$ is not an element of $\operatorname{\mathcal{S}}$. This is explicitly allowed because logical operators are not necessarily observables and hence don’t have to adhere to parity superselection rules. They are therefore allowed to flip the parity of a logical state similarly to (76). From a physical perspective, this can be interpreted for example as the logical quantum states interacting with each other through the exchange of an odd number of fermions. In the existing literature this is referred to as quasiparticle poisoning [12, 15].
IV.4 Stabilizers from p-Cliffords
As already implicitly shown in (64), any $[[n,0]]_{f}$ stabilizer basis can be generated by applying a single p-Clifford $\operatorname{\gamma}\in\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ to each of the standard basis strings $\operatorname{\mu}(b_{i})$ of $M_{\text{std}}$ in (65). However, things become non-trivial when considering stabilizers with $k>0$, as those might or might not contain the element ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$. And since $j$ gets mapped to itself under application of any orthogonal matrix (as a direct consequence of such matrices preserving parity), applying a p-Clifford to any stabilizer (not) containing ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$ can only result in some other stabilizer (not) containing ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$. This is problematic e.g. if one wants to be able to prepare an arbitrary stabilizer using an encoding quantum circuit, or if one wants to uniformly sample over all possible stabilizers numerically.
There are multiple ways to approach this issue, the one we describe here is inspired by [15] and uses the addition of two unoccupied ancilla modes to the stabilizer $\operatorname{\mathcal{S}}$ to remove any potential presence of ±$\pm\operatorname{\bar{\operatorname{\mu}}}(j)$. One can then use (even-parity) braiding operators (58) to diagonalize the given basis of $\operatorname{\mathcal{S}}$ such that it is represented by the first $r$ rows of (65), each with two additional entries representing the ancilla appended at the end. During the process of diagonalization these ancilla modes might be occupied, but one can always choose the braiding operators such that they won’t be in the end. The composition of the inverse braiding operations (with additional phase changes) hence provides an encoding p-Clifford for $\operatorname{\mathcal{S}}$ on the extended Hilbert space.
The following algorithm performs this computation in the symplectic representation, i.e. given an ordered basis $(b_{1},\ldots,b_{r})$ of an isotropic subspace $M\subset\mathbb{F}_{2}^{2n,+}$ that does not contain $j$ [^9] We assume that ancilla modes were introduced in advance., the routine returns an orthogonal matrix $S\in\operatorname{O}(2n)$ that maps $M^{r}_{\text{std}}$ (the isotropic subspace spanned by the first $r$ rows of (65)) to $M$, and in particular the standard basis of $M^{r}_{\text{std}}$ to $\{b_{1},\ldots,b_{r}\}$.
Algorithm 1 STABCLIFFORD($M$)
$M$ as basis $(b_{1},\ldots,b_{r})$, $j\notin M$
$S\in\operatorname{O}(2n)$ s.th. $SM^{r}_{\text{std}}=M$
$S\leftarrow I\in\operatorname{O}(2n)$
for $i=1,\ldots,r$ do
     $m\leftarrow S\,b_{i}$
     $m^{\prime}\leftarrow m_{(2i-1),\ldots,2n}\in\mathbb{F}_{2}^{2(n-i+1),+}$
     $e^{\prime}\leftarrow(1,1,0,\ldots,0)\in\mathbb{F}_{2}^{2(n-i+1),+}$
     if $m^{\prime}=e^{\prime}$ then
         $a^{\prime},b^{\prime}\leftarrow(0,1,1,0,\ldots,0)^{T}\in\mathbb{F}_{2}^{2(n-i+1),+}$
     else
         $a^{\prime},b^{\prime}\leftarrow\text{HOUSEHOLDER}(e^{\prime},m^{\prime})$
     end if
     $a\leftarrow m_{1,\ldots,(2i-2)}\oplus a^{\prime}\in\mathbb{F}_{2}^{2n,+}$
     $b\leftarrow(0,\ldots,0)^{T}\oplus b^{\prime}\in\mathbb{F}_{2}^{2n,+}$
     $S\leftarrow h_{b}h_{a}\,S$
end for
return $S^{T}$
We prove that this algorithm works using induction with regard to the iteration steps:
  1. 1.
    Starting with the first basis element $S\,b_{1}=b_{1}$, because of our initial requirements we can always find a set of Householder reflections $h_{b}h_{a}$ that map it to the standard basis vector $(1,1,\ldots,0)^{T}$ (see Appendix B). We therefore add them to $S$ such that they are applied first.
  2. 2.
    Assuming that $S$ is composed of reflections that map the first $k$ basis elements to their respective standard basis elements, we now have to find a pair of reflections that map $m=S\,b_{k+1}$ to the next standard basis element.
  3. 3.
    Since $m$ must commute with all previous elements, we know that the first $22k-$ 2 entries of $m$ are composed of consecutive pairs of either ones or zeroes. We also know that the remaining entries can’t be all zeros or all ones, because then $m$ would either not be linearly independent, or one would have $j\in\operatorname{span}(S\,b_{1},\ldots,S\,b_{r})$ and therefore also $j\in\operatorname{span}(b_{1},\ldots,b_{r})$, both of which go against the initial assumptions.
  4. 4.
    Let $m^{\prime}$ be the (even-parity) vector corresponding to those $2(n-i+1)$ trailing entries. We want to find non-identity Householder reflections $h_{a^{\prime}},h_{b^{\prime}}$ that map $(1,1,0,\ldots,0)$ to $m^{\prime}$, which as mentioned before is always possible.
  5. 5.
    Using the corresponding even-parity vectors $a^{\prime}\neq 0$ and $b^{\prime}$, we can construct even-parity vectors $a$ and $b$ by adding the first $22k-$ 2 entries of $m$ to the front of $a^{\prime}$, and adding that many zeros to the front of $b^{\prime}$. It is straightforward to check that the corresponding reflections $h_{a},h_{b}$ map $m$ to the $(k+1)$th standard basis element, while keeping the previous $k$ elements invariant. One can therefore update $S$ and proceed to the next iteration.
  6. 6.
    Finally, the inverse/transpose of $S$ is returned.
Since Householder reflections representing the braiding operators can be found and applied to vectors in $\mathcal{O}(n)$ time (see Appendix B), the total complexity of this algorithm is $\mathcal{O}(n^{2}r)$. In [15], a variation of this procedure is encoded as a quantum circuit, which uses the generating set of braiding operators (see Theorem 2) to perform Gaussian elimination.
V p-Cliffords as a Unitary Design
It is a well-known fact that the Clifford group $\operatorname{\mathcal{C}}_{2n}$ (in the fundamental representation) forms a 3-design of the unitary group $\operatorname{U}(2^{n})$ [29].
Definition 11 (Unitary $t$-Design).
An ensemble $\operatorname{\mathcal{E}}$ of unitary operators acting on a $N$-dimensional Hilbert space $\operatorname{\mathcal{H}}$ is said to be a $t$-design iff the ensemble average replicates the unitary Haar average:
[…]
for any linear operator $O$ acting on $\operatorname{\mathcal{H}}^{\otimes t}$.
Equivalently, an unitary ensemble $\operatorname{\mathcal{E}}$ is a $t$-design iff the associated $t$-th frame potential
$\operatorname{\mathcal{F}}_{t}(\operatorname{\mathcal{E}})\coloneqq\operatorname*{{\mathbb{E}}}_{U,V\in\operatorname{\mathcal{E}}}\,\lvert\operatorname{tr}(UV^{\dagger})\rvert^{2t}$ (78)
matches the $t$-th frame potential $\operatorname{\mathcal{F}}_{t}(\operatorname{U}(N))\equiv\upsilon(t,N)$ of the unitary group in $N$ dimensions. The latter can be shown to have the following relevant properties [29]:
$\upsilon(t,N)=\begin{cases}\frac{(2t)!}{t!(t+1)!},&N=2\\ t!,&N\geq t\end{cases}.$ (79)
Per definition, any ensemble that is a $t$-design ($t>1$) is also a $(t-1)$-design. Additionally, if the ensemble forms a group $\mathcal{G}$, $\operatorname{\mathcal{F}}_{t}(\mathcal{G})$ simplifies to
$\operatorname{\mathcal{F}}_{t}(\mathcal{G})=\operatorname*{{\mathbb{E}}}_{U\in\mathcal{G}}\,\lvert\operatorname{tr}U\rvert^{2t}.$ (80)
As per Lemma 2 of [29], for any subgroup $\mathcal{G}\leq\operatorname{\mathcal{C}}_{2n}$ of the Clifford group satisfying $\mathcal{G}\geq\operatorname{\mathcal{M}}_{2n}$ (including $\operatorname{\mathcal{C}}_{2n}$ itself), the frame potential can be expressed in terms of the subgroup’s symplectic representation $G\leq\operatorname{Sp}(2n)$:
$\operatorname{\mathcal{F}}_{t}(\mathcal{G})=\frac{1}{|G|}\sum_{S\in G}f(S)^{t-1},$ (81)
where $f(S)$ denotes the number of fixed points (i.e. unit eigenvectors) of $S$ in $\mathbb{F}_{2}^{2n}$. This is in part because the absolute trace $|\operatorname{tr}(\operatorname{\mu}(a)\gamma(S))|^{2}$ of a general Clifford is either zero or equal to the number of Majorana strings it commutes with.
According to the orbit-stabilizer relation, $\operatorname{\mathcal{F}}_{t}(\mathcal{G})$ is also equivalent to the number of orbits of $G$ on $(\mathbb{F}_{2}^{2n})^{\times(t-1)}$. $\mathcal{G}$ is therefore a 2-design iff $G$ has 2 orbits on $\mathbb{F}_{2}^{2n}$ (and is transitive on $\mathbb{F}_{2}^{2n}\setminus\{0\}$). It is also a 3-design iff $G$ has 6 orbits on $(\mathbb{F}_{2}^{2n})^{\times 2}$ (or 5 if $n=1$), and so on. Using these last statements, it is easy to show that the p-Clifford subgroup is not a 2-design on $\operatorname{\mathcal{H}}$. This is because $\operatorname{O}(2n)$ has 4 orbits on $\mathbb{F}_{2}^{2n}$, corresponding to
$\{0\},\quad\{j\},\quad\mathbb{F}_{2}^{2n,+}\setminus\{0,j\},\quad\mathbb{F}_{2}^{2n,-}.$
V.1 Parity-Restricted Frame Potential
However, based on the parity superselection rules we can argue that this is not the correct way to think about $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ as a $t$-design. The frame potential we are interested in should only care about how p-Cliffords (or some other ensemble) act on a (maximal) subspace of $\operatorname{\mathcal{H}}$ with fixed parity. Without loss of generality, we choose the even-parity sector for our consideration and hence define:
Definition 12 (Parity-Restricted Frame Potential).
Let $\operatorname{\mathcal{E}}$ be an unitary ensemble acting on a Hilbert space $\operatorname{\mathcal{H}}$ such that it preserves fermion parity, i.e. $U^{\dagger}(-1)^{F}U=(-1)^{F}$ for all $U\in\operatorname{\mathcal{E}}$. Then the corresponding parity-restricted frame potential is defined as
$\operatorname{\mathcal{F}}^{+}_{t}(\operatorname{\mathcal{E}})\coloneqq\operatorname*{{\mathbb{E}}}_{U,V\in\operatorname{\mathcal{E}}}\,\lvert\operatorname{tr}(\Pi_{+}UV^{\dagger})\rvert^{2t},$ (82)
where $\Pi_{+}$ projects onto the even-parity subspace of $\operatorname{\mathcal{H}}$:
$\Pi_{+}\equiv\frac{1}{2}\left(I+(-1)^{F}\right).$ (83)
Any unitary ensemble $\operatorname{\mathcal{E}}$ acting on a Hilbert space of dimension $2^{n-1}=\operatorname{tr}(\Pi_{+})$ can be extended to a parity-preserving ensemble $\overline{\operatorname{\mathcal{E}}}$ in $2^{n}$ dimensions by having its elements act trivially on the 2-dimensional odd-parity subspace. In this case the parity-restricted frame potential of $\overline{\operatorname{\mathcal{E}}}$ is the same as the ordinary one for $\operatorname{\mathcal{E}}$:
$\operatorname{\mathcal{F}}^{+}_{t}(\overline{\operatorname{\mathcal{E}}})=\operatorname{\mathcal{F}}_{t}(\operatorname{\mathcal{E}}).$ (84)
The same approach is however not feasible for $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$, as the sub-ensemble acting on the even-parity subspace is not known a priori. Instead, one can prove the following:
Theorem 4.
For any subgroup $\operatorname{\mathcal{M}}^{+}_{2n}\leq\mathcal{G}\leq\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ of the p-Clifford group (including $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ itself), the parity-restricted frame potential $\operatorname{\mathcal{F}}^{+}_{t}(\mathcal{G})$ can be expressed in terms of the subgroup’s orthogonal representation $G\leq\operatorname{O}(2n)$:
$\operatorname{\mathcal{F}}^{+}_{t}(\mathcal{G})=\frac{1}{|G|}\sum_{S\in G}\left[\frac{f_{+}(S)+c_{+}(S)}{2}\right]^{t-1},$ (85)
where $f_{+}(S)$ is the number of even-parity fixed points of $S$ (satisfying $Sv=v$) and $c_{+}(S)$ is the number of even-parity complemented points (satisfying $Sv=v^{c}$).
A formal proof is provided in Appendix A.2, but for now the expression $(f_{+}(S)+c_{+}(S))/2$ being summed over on the right side of (85) has the intuitive interpretation of being the number of fixed points in the quotient vector space $\mathbb{F}_{2}^{2n,+}/\sim^{c}$ associated to the equivalence relation $v\sim^{c}v^{c}=v+j$. This is because the parity-check string $\operatorname{\mu}(j)\equiv(-1)^{F}$ commutes with all elements of the even-parity subgroup $\operatorname{\mathcal{M}}^{+}_{2n}$ and must therefore be equivalent to the identity $I\equiv\operatorname{\mu}(0)$ due to Schur’s lemma. This identification together with the composition property (6) then leads to the aforementioned equivalence relation and the reason why we also have to consider $c_{+}(S)$. The additional factor of 1/2 ensures that each pair $(v,v^{c})$ is only counted once, since if $v$ is a fixed/complemented point of $S$, then so is $v^{c}$.
Finally, the number of complemented points $c_{+}(S)$ is either 0 or equal to the number of fixed points $f_{+}(S)$. One can see this from the fact that the difference of two complemented points is always a fixed point, and therefore a single complemented point suffices to construct $f_{+}(S)$ distinct such points.
Using the orbit-stabilizer relation we can therefore overall conclude that $\operatorname{\mathcal{F}}^{+}_{t}(\mathcal{G})$ for any subgroup $\operatorname{\mathcal{M}}^{+}_{2n}\leq\mathcal{G}\leq\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ is equal to the number of orbits of $G\leq\operatorname{O}(2n)$ on $(\mathbb{F}_{2}^{2n,+}/\sim^{c})^{\times(t-1)}$.
V.2 Closing the Circle
One can now compare the parity-restricted frame potentials of $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ and $\operatorname{\mathcal{C}}_{2n-2}\otimes I_{-}$, to see if the former does a better job at replicating $\upsilon(t,2^{n-1})$ (79). This comparison can be done exactly for small $t$ and $n$ (e.g. by evaluating the orbits), but for larger values a Monte Carlo-based numerical analysis is more suitable.
What we found (exactly for $n=2,3,4$ and numerically for $n\geq 4$) is that $\operatorname{\mathcal{F}}^{+}_{t}(\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n})$ does match $\operatorname{\mathcal{F}}^{+}_{t}(\operatorname{\mathcal{C}}_{2n}\otimes I_{-})=\operatorname{\mathcal{F}}_{t}(\operatorname{\mathcal{C}}_{2n})$ precisely. Using these results we were then able to formulate an exact proof:
Theorem 5.
The parity-restricted frame potential of $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ is equal to the ordinary frame potential of $\operatorname{\mathcal{C}}_{2n-2}$:
$\operatorname{\mathcal{F}}^{+}_{t}(\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n})=\operatorname{\mathcal{F}}_{t}(\operatorname{\mathcal{C}}_{2n-2})\quad\forall n\geq 2,t\geq 1$ (86)
Proof.
Using (81) and (85), it is sufficient to show that the effective action of $\operatorname{O}(2n)$ on $\mathbb{F}_{2}^{2n,+}/\sim^{c}$ is equal to ($2^{2n-1}$ copies of) $\operatorname{Sp}(2n-2)$, which is done in Appendix A.3. This then implies that the distributions of fixed points $f(S)$ ($S\in\operatorname{Sp}(2n)$) and $(f_{+}(S)+c_{+}(S))/2$ ($S\in\operatorname{O}(2n)$) are the same up to an overall factor of $2^{2n-1}$, which cancels with part of the denominator in (85):
$|\operatorname{O}(2n)|=2^{2n-1}\cdot|\operatorname{Sp}(2n-2)|.$ (87)
So while this implies that the p-Clifford group $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ is also a 3-design on a fixed-parity subspace for all $n\geq 2$, there seems to be no obvious advantage compared to the naive ensemble $\operatorname{\mathcal{C}}_{2n-2}\otimes I_{-}$, besides the former being the more natural choice due to not requiring an explicit Hilbert space decomposition.
VI Conclusions
In this paper, we have compiled a comprehensive overview of Majorana strings and the corresponding Clifford operators through the lens of their symplectic representation. Assuming parity superselection, we could identify a subgroup of parity-preserving Cliffords, which is represented by binary orthogonal matrices. We also showed how these p-Cliffords can be efficiently generated from a small set of braiding operators, and how they can be used to prepare any even-parity stabilizer code.
A frame potential analysis of the p-Clifford subgroup restricted to a fixed-parity sector of the Hilbert space showed that it is equivalent to the ordinary Clifford group acting on the same subspace. The p-Clifford group is thus also a 3-design.
One application of the methods developed in this paper is the study of fermionic random tensor network models of SYK as envisioned in [21] and perhaps other fermionic tensor network models of quantum gravity [30]. The results presented here might also be useful in analyzing other Majorana-based systems, such as in fermionic quantum codes and fermionic random circuits, and could perhaps be combined with fermionic tensor network machinery [31]. It could also be interesting to investigate fermionic analogs of the Wigner function, which is a useful tool for characterizing qudit stabilizer state [23], although there are known difficulties in the qubit case which may be inherited in the fermionic context.
VII Acknowledgments
V.B. would like to thank Nadie LiTenn and Shiyong Guo for some insightful discussions and comments. We acknowledge support from the AFOSR under grant number FA9550-19-1-0360 (V.B. and B.G.S.).
Appendix A Proofs
A.1 Theorem 1
Proof.
  1. 1.
    Using the fact that $\operatorname{\gamma}$ is unitary, one can rewrite (13) into the equivalent statement
    $\begin{split}&[\operatorname{\gamma}\operatorname{\mu}(v)\operatorname{\gamma}^{\dagger}][\operatorname{\gamma}\operatorname{\mu}(v^{\prime})\operatorname{\gamma}^{\dagger}]\\ ={}&(-1)^{\left\langle v,v^{\prime}\right\rangle}[\operatorname{\gamma}\operatorname{\mu}(v^{\prime})\operatorname{\gamma}^{\dagger}][\operatorname{\gamma}\operatorname{\mu}(v)\operatorname{\gamma}^{\dagger}]\\ \Longleftrightarrow\quad\quad&\operatorname{\mu}(S(v))\operatorname{\mu}(S(v^{\prime}))\\ ={}&(-1)^{\left\langle v,v^{\prime}\right\rangle}\operatorname{\mu}(S(v^{\prime}))\operatorname{\mu}(S(v)),\end{split}$ (88)
    where we used the fact that $c(v)\neq 0$. But since (13) must also hold when we directly set $v\rightarrow S(v)$, we get the following requirement:
    $\left\langle S(v),S(v^{\prime})\right\rangle\stackrel{{\scriptstyle!}}{{=}}\left\langle v,v^{\prime}\right\rangle\quad\forall v,v^{\prime}\in\mathbb{F}_{2}^{2n}.$ (89)
[…]
    $\begin{split}&\left\langle S(v+v^{\prime}),S(v^{\prime\prime})\right\rangle\\ ={}&\left\langle v+v^{\prime},v^{\prime\prime}\right\rangle\\ ={}&\left\langle v,v^{\prime\prime}\right\rangle+\left\langle v^{\prime},v^{\prime\prime}\right\rangle\\ ={}&\left\langle S(v),S(v^{\prime\prime})\right\rangle+\left\langle S(v^{\prime}),S(v^{\prime\prime})\right\rangle\\ ={}&\left\langle S(v)+S(v^{\prime}),S(v^{\prime\prime})\right\rangle\quad\forall v,v^{\prime},v^{\prime\prime}\in\mathbb{F}_{2}^{2n}.\end{split}$ (90)
    With $S$ being linear it then immediately follows from Definition 6 that it must also be symplectic.
    The converse direction follows from the fact that both $\operatorname{\mu}(v)$ and $\operatorname{\mu}(Sv)$ define irreducible representations of $\mathbb{F}_{2}^{2n}$ because of Lemma 2 and $S$ being linear:
    $\frac{1}{|\mathbb{F}_{2}^{2n}|}\sum_{v\in\mathbb{F}_{2}^{2n}}|\operatorname{tr}(\operatorname{\mu}(Sv))|^{2}=1.$ (91)
    Both representations are therefore equivalent and hence (up to an overall sign) connected by a unitary transformation $\gamma$ that induces $S$.
  2. 2.
    Assuming $\operatorname{\gamma}(S)$ satisfies (38), then $\operatorname{\mu}(a)\operatorname{\gamma}(S)$ with $a\in\mathbb{F}_{2}^{2n}$ does so as well:
    $\begin{split}&[\operatorname{\mu}(a)\operatorname{\gamma}(S)]\operatorname{\mu}(v)[\operatorname{\mu}(a)\operatorname{\gamma}(S)]^{\dagger}\\ ={}&\operatorname{\mu}(a)\operatorname{\gamma}(S)\operatorname{\mu}(v)\operatorname{\gamma}^{\dagger}(S)\operatorname{\mu}(a)\\ ={}&c(v)\cdot\operatorname{\mu}(a)\operatorname{\mu}(Sv)\operatorname{\mu}(a)\\ ={}&(-1)^{\left\langle Sv,a\right\rangle}c(v)\cdot\operatorname{\mu}(Sv)\\ \equiv{}&c^{\prime}(v)\cdot\operatorname{\mu}(Sv),\end{split}$ (92)
    where we used (13) and the fact that $\operatorname{\mu}(a)^{2}=I$.
    Conversely, let $\operatorname{\gamma}(S)$ and $\operatorname{\gamma}^{\prime}(S)$ be two Cliffords with the same symplectic action. Then dividing both cases of (46) yields
    $\left[\frac{c(v)}{c^{\prime}(v)}\right]\left[\frac{c(v^{\prime})}{c^{\prime}(v^{\prime})}\right]=\frac{c(v+v^{\prime})}{c^{\prime}(v+v^{\prime})}$ (93)
    for all $v,v^{\prime}\in\mathbb{F}_{2}^{2n}$. This together with $c(v)^{2}=c(0)=1$ (and the symplectic form being non-degenerate) implies that one can write without loss of generality
    $\frac{c(v)}{c^{\prime}(v)}=(-1)^{\left\langle v,a^{\prime}\right\rangle}=(-1)^{\left\langle Sv,a\right\rangle},$ (94)
    where we set $a^{\prime}\equiv S^{-1}a\in\mathbb{F}_{2}^{2n}$. This coincides with the form of the phase induced by an additional Majorana string $\operatorname{\mu}(a)$, as shown in (LABEL:eq:cliff_majo_action).
  3. 3.
    $\operatorname{\mathcal{C}}_{2n}/\sim$ forms a representation of $\operatorname{Sp}(2n)$, because:
    $\begin{split}&\operatorname{\gamma}(S)\operatorname{\gamma}(T)\operatorname{\mu}(v)\operatorname{\gamma}^{\dagger}(T)\operatorname{\gamma}^{\dagger}(S)\\ \propto{}&\operatorname{\gamma}(S)\operatorname{\mu}(Tv)\operatorname{\gamma}^{\dagger}(S)\\ \propto{}&\operatorname{\mu}(STv)\\ \propto{}&\operatorname{\gamma}(ST)\operatorname{\mu}(v)\operatorname{\gamma}^{\dagger}(ST)\quad\forall v\in\mathbb{F}_{2}^{2n}.\end{split}$ (95)
    which implies that $\operatorname{\gamma}(S)\operatorname{\gamma}(T)\sim\operatorname{\gamma}(ST)$ and therefore $[\operatorname{\gamma}(S)][\operatorname{\gamma}(T)]=[\operatorname{\gamma}(ST)]$.
A.2 Theorem 4
Proof.
Because $\mathcal{G}\leq\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ forms a group and therefore satisfies $UV^{\dagger}\in\mathcal{G}$ for all $U,V\in\mathcal{G}$, the frame potential can be simplified to be
$\begin{split}\operatorname{\mathcal{F}}^{+}_{t}(\mathcal{G})&=\operatorname*{{\mathbb{E}}}_{U\in\mathcal{G}}\,\lvert\operatorname{tr}(\Pi_{+}U)\rvert^{2t}\\ &=\frac{1}{\lvert\mathbb{F}_{2}^{2n,+}\rvert}\frac{1}{|G|}\sum_{S\in G}\sum_{a\in\mathbb{F}_{2}^{2n,+}}\lvert\operatorname{tr}[\Pi_{+}\operatorname{\gamma}(S)\operatorname{\mu}(a)]\rvert^{2t}.\end{split}$ (96)
We first show that $\lvert\operatorname{tr}(\Pi_{+}U)\rvert^{2}$ with $U=\operatorname{\mu}(a)\operatorname{\gamma}(S)\in\mathcal{G}$ is either zero or equal to $(f_{+}(S)+c_{+}(S))/2$. For that we evaluate the expression as a sum over the Bell basis $\lvert\Phi_{n}(v)\rangle\equiv(\operatorname{\mu}(v)\otimes I)\lvert\Phi_{n}^{+}\rangle$ of $\operatorname{\mathcal{H}}\otimes\operatorname{\mathcal{H}}$, where $\lvert\Phi_{n}^{+}\rangle\equiv\lvert\Phi^{+}\rangle^{\otimes n}$ denotes $n$ copies of canonical Bell state:
$\begin{split}&\lvert\operatorname{tr}(\Pi_{+}U)\rvert^{2}\\ ={}&\operatorname{tr}(\Pi_{+}U\otimes\Pi_{+}U^{\dagger})\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n}}\langle\Phi_{n}(v)\rvert\Pi_{+}U\otimes\Pi_{+}U^{\dagger}\lvert\Phi_{n}(v)\rangle\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n}}\langle\Phi_{n}^{+}\rvert\operatorname{\mu}(v)\Pi_{+}U\operatorname{\mu}(v)\otimes\Pi_{+}U^{\dagger}\lvert\Phi_{n}^{+}\rangle\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n}}\langle\Phi_{n}^{+}\rvert\operatorname{\mu}(v)\Pi_{+}U\operatorname{\mu}(v)U^{\dagger}\Pi_{+}\otimes I\lvert\Phi_{n}^{+}\rangle\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n}}c(v)\,\langle\Phi_{n}^{+}\rvert\operatorname{\mu}(v)\Pi_{+}\operatorname{\mu}(Sv)\Pi_{+}\otimes I\lvert\Phi_{n}^{+}\rangle\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n,+}}c(v)\,\langle\Phi_{n}^{+}\rvert\operatorname{\mu}(v)\operatorname{\mu}(Sv)\Pi_{+}\otimes I\lvert\Phi_{n}^{+}\rangle\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n,+}}\frac{c(v)}{2^{n}}\,\operatorname{tr}[\operatorname{\mu}(v)\operatorname{\mu}(Sv)\Pi_{+}]\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n,+}}\frac{c(v)}{2^{n+1}}\operatorname{tr}[\operatorname{\mu}(v)\operatorname{\mu}(Sv)+\operatorname{\mu}(v)\operatorname{\mu}(Sv)\operatorname{\mu}(j)]\\ ={}&\sum_{v\in\mathbb{F}_{2}^{2n,+}}\frac{c(v)}{2}(\delta_{Sv,v}+\zeta(j,v)\,\delta_{Sv,v^{c}})\\ \end{split}$ (97)
We can now analyze the structure of $c(v)$ for both cases:
  • In the case of $Sv=v$, it follows directly from (46) that the coefficients of any two (even-parity) fixed points $v,v^{\prime}$ satisfy $c(v)\,c(v^{\prime})=c(v+v^{\prime})$. This together with $c(v)^{2}=1$ implies that we can choose $c(v)=(-1)^{a^{T}v}$ without loss of generality and in accordance with (LABEL:eq:cliff_majo_action).
  • In the case of $Sv=v^{c}$, for any two (even-parity) complemented points $v,v^{\prime}$ (46) can we rewritten as
    $c(v)\,c(v^{\prime})=\frac{\zeta(v,v^{\prime})}{\zeta(v^{c},(v^{\prime})^{c})}\cdot c(v+v^{\prime}).$ (98)
    But since the sum of two complemented points is always a fixed point, we can identify $c(v+v^{\prime})=(-1)^{a^{T}(v+v^{\prime})}$, as before. Additionally, by evaluating $\operatorname{\mu}(j)\operatorname{\mu}(v)\operatorname{\mu}(v^{\prime})\operatorname{\mu}(j)$ in two different orders one arrives at
    $\zeta(v,v^{\prime})=\zeta(j,v)\,\zeta(j,v^{\prime})\,\zeta(v^{c},(v^{\prime})^{c}).$ (99)
    Plugging this all into (98) and using $\zeta(j,v)^{2}=1$ (since $\left\langle j,v\right\rangle=0$), results in
    $\left[c(v)\,\zeta(j,v)\right]\left[c(v^{\prime})\,\zeta(j,v^{\prime})\right]=(-1)^{a^{T}(v+v^{\prime})},$ (100)
    which therefore implies $c(v)\,\zeta(j,v)=(-1)^{a^{T}v}$.
Applying these results to the evaluation of $\lvert\operatorname{tr}(\Pi_{+}U)\rvert^{2}$ then leaves us with
$\begin{split}\lvert\operatorname{tr}(\Pi_{+}U)\rvert^{2}&=\sum_{v\in\mathbb{F}_{2}^{2n,+}}\frac{(-1)^{a^{T}v}}{2}(\delta_{Sv,v}+\delta_{Sv,v^{c}})\\ &=\frac{1}{2}\left(f_{+}(S)+c_{+}(S)\right)\quad(\text{or }0),\end{split}$ (101)
where the second line is a direct consequence of
$\sum_{v\in M}(-1)^{a^{T}v}=|M|\,\delta_{a\in M^{\perp}}$ (102)
for any subspace $M\subseteq\mathbb{F}_{2}^{2n}$ and $M^{\perp}=\{v\in\mathbb{F}_{2}^{2n}\,|\,v^{T}v^{\prime}=0\quad\forall v^{\prime}\in M\}$. This also implies that
$\sum_{a\in\mathbb{F}_{2}^{2n,+}}\lvert\operatorname{tr}[\Pi_{+}\operatorname{\gamma}(S)\operatorname{\mu}(a)]\rvert^{2}=2^{2n-1},$ (103)
since $(\mathbb{F}_{2}^{2n,+})^{\perp}\cap\mathbb{F}_{2}^{2n,+}=\{0,j\}$, and both elements are fixed points. Hence in this sum $2^{2n}/(f_{+}(S)+c_{+}(S))$ terms are non-zero and therefore equal to (101). It follows that
$\begin{split}&\sum_{a\in\mathbb{F}_{2}^{2n,+}}\lvert\operatorname{tr}[\Pi_{+}\operatorname{\gamma}(S)\operatorname{\mu}(a)]\rvert^{2t}\\ ={}&\frac{2^{2n}}{(f_{+}(S)+c_{+}(S))}\cdot\left[\frac{f_{+}(S)+c_{+}(S)}{2}\right]^{t}\\ ={}&\lvert\mathbb{F}_{2}^{2n,+}\rvert\cdot\left[\frac{f_{+}(S)+c_{+}(S)}{2}\right]^{t-1},\end{split}$ (104)
and finally
$\operatorname{\mathcal{F}}^{+}_{t}(\mathcal{G})=\ \frac{1}{|G|}\sum_{S\in G}\left[\frac{f_{+}(S)+c_{+}(S)}{2}\right]^{t-1}.$ (105)
A.3 Lemma for Theorem 5
Lemma 4.
The orthogonal group $\operatorname{O}(2n)$ restricted to $\mathbb{F}_{2}^{2n,+}/\sim^{c}$ (where $v\sim^{c}v^{c}=v+j$) is equal to ($2^{2n-1}$ copies of) $\operatorname{Sp}(2n-2)$, and therefore $\operatorname{Sp}(2n-2)\leq\operatorname{O}(2n)$.
Proof.
Without loss of generality, we start with the following basis of $\mathbb{F}_{2}^{2n,+}$ as a $(2n-1)\times 2n$ matrix, with its elements as row vectors:
$B=\begin{pmatrix}1&0&1&0&0&\cdots&0&0&0&0\\ 1&1&0&0&0&\cdots&0&0&0&0\\ 1&1&1&0&1&\cdots&0&0&0&0\\ 1&1&1&1&0&\cdots&0&0&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\vdots\\ 1&1&1&1&1&\cdots&1&0&1&0\\ 1&1&1&1&1&\cdots&1&1&0&0\\ \hline\cr 1&1&1&1&1&\cdots&1&1&1&1\end{pmatrix}.$ (106)
Going to $\mathbb{F}_{2}^{2n,+}/\sim^{c}$ is then achieved by removing the last line of $B$ (corresponding to $j$) and identifying all other lines with their respective complement [^10] Identifying each basis element with its complement is only really necessary to ensure the basis is still complete, it does not affect anything else.. We will refer to this new basis as $\widetilde{B}$, and any element of $\mathbb{F}_{2}^{2n,+}/\sim^{c}$ can be expressed as $\widetilde{B}^{T}v$ with $v\in\mathbb{F}_{2}^{2n-2}$. The canonical inner product of any two elements then is
$\left(\widetilde{B}^{T}v\right)^{T}\left(\widetilde{B}^{T}v^{\prime}\right)=v^{T}\left(\widetilde{B}\widetilde{B}^{T}\right)v^{\prime}=v^{T}\operatorname{\eta}v^{\prime},$ (107)
where $\operatorname{\eta}$ is the symplectic form in the Pauli basis (25). For any orthogonal matrix $S\in\operatorname{O}(2n)$ there therefore must be a symplectic matrix $\widetilde{S}\in\operatorname{Sp}(2n-2)$ such that
$\widetilde{B}S^{T}=\widetilde{S}^{T}\widetilde{B}.$ (108)
Conversely, for any $\widetilde{S}\in\operatorname{Sp}(2n-2)$ there are $2^{2n-1}$ orthogonal matrices $S\in\operatorname{O}(2n)$ with that effective action on $\mathbb{F}_{2}^{2n,+}/\sim^{c}$, because they can only differ by how they map the one-dimensional odd-parity subspace to one of the $2^{2n-1}$ possible options. ∎
Appendix B Sampling p-Cliffords
In this appendix we provide an efficient sampling algorithm for elements of the p-Clifford group $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ in terms of their orthogonal representation $S\in\operatorname{O}(2n)$. It is inspired by the more general algorithm for sampling arbitrary symplectics provided in [27].
Since orthogonal matrices exist for any even or odd dimension $N\geq 1$, the algorithm can be defined recursively to sample all of them. It uses the fact that any two vectors $v,w\in\mathbb{F}_{2}^{N}\setminus\{j,0\}$ of the same parity are connected by either one or two Householder reflections $h_{1},h_{2}\in\operatorname{O}(N)$, such that
$h_{2}h_{1}v=w,$ (109)
where $h_{2}$ is potentially just the identity.
Algorithm 2 HOUSEHOLDER($v,w$)
$v,w\in\mathbb{F}_{2}^{N},\,\operatorname{p}(v)=\operatorname{p}(w),\,v\neq j\neq w$
$a,b\in\mathbb{F}_{2}^{N,+},\,h_{b}h_{a}v=w$
if $v^{T}w=1-\operatorname{p}(v)$ or $v=w$ then
     $a\leftarrow w-v$
     $b\leftarrow 0$
     return
end if
$a\leftarrow 0$
for $i=1,\ldots,N$ do
     if $v_{i}=0$ and $w_{i}=0$ then
         for $i^{\prime}=1,\ldots,N$ do
              if $v_{i^{\prime}}=1$ and $w_{i^{\prime}}=1$ then
                  $a_{i}\leftarrow 1,\,a_{i^{\prime}}\leftarrow 1$
                  $b\leftarrow w-v-a$
                  return
              end if
         end for
         break
     end if
end for
for $i=1,\ldots,N$ do
     if $v_{i}=1$ and $w_{i}=0$ then
         $a_{i}\leftarrow 1$
         break
     end if
end for
for $i=1,\ldots,N$ do
     if $v_{i}=0$ and $w_{i}=1$ then
         $a_{i}\leftarrow 1$
         break
     end if
end for
$b\leftarrow w-v-a$
return
This algorithm has $\mathcal{O}(N)$ runtime and returns the even-parity vectors defining the Householder reflections connecting $v$ and $w$. It acts as follows:
  1. 1.
    We first check if $v^{T}w=1-\operatorname{p}(v)$ (or $v=w$), because then $v^{T}(w-v)=1$ (or $v^{T}(w-v)=0$ if $v=w$ since nothing has to be done), and hence the Householder $h_{w-v}$ satisfies the desired property.
  2. 2.
    Next we check if $v$ and $w$ have a common index $i$ for which $v_{i}=0=w_{i}$. If such an index exists, we also check if $v$ and $w$ share an index $i^{\prime}$ for which $v_{i^{\prime}}=1=w_{i^{\prime}}$. If both indices exist, then $a$ with $a_{i}=1=a_{i^{\prime}}$ as the only non-zero elements satisfies $a^{T}v=1=a^{T}w$ and thus together with $b=w-v-a$ provides a valid pair or reflections.
  3. 3.
    We can now assume that $v$ is not fully contained in $w$, or vice versa. If that were the case, they would have a common zero since neither of them can be equal to $j$. Therefore the only other option is that $v_{i}=1,w_{i}=0$ for (at least) one index $i$, and $v_{i^{\prime}}=0,w_{i^{\prime}}=1$ for another index $i^{\prime}$. The vector $a$ that only has ones at $i$ and $i^{\prime}$, and zeroes everywhere else, therefore describes a valid reflection. The other one is again described by $b=w-v-a$.
The actual routine for sampling orthogonal matrices is now straightforward.
Algorithm 3 ORTHOGONAL($N,i$)
$1\leq i\leq|\operatorname{O}(N)|$
$g_{i}\in\operatorname{O}(N),\quad g_{i}\neq g_{i^{\prime}}\Leftrightarrow i\neq i^{\prime}$
if $N=1$ then
     $g_{i}\leftarrow 1$
     return
end if
$p\leftarrow 2^{N-1}$triangleright Number of odd-parity vectors of length $N$
$i_{p}\leftarrow i-1\mod p$
$i_{r}\leftarrow(i-1)/p+1$triangleright Index for next recursive step.
$b_{i}\leftarrow\text{binary representation of }i_{p}\text{ as vector in }\mathbb{F}_{2}^{N-1}$
$e\leftarrow(1,0,\ldots,0)^{T}\in\mathbb{F}_{2}^{N}$
$f\leftarrow(\operatorname{p}(b_{i})+1)\oplus b_{i}\in\mathbb{F}_{2}^{N}$
$a,b\leftarrow\text{HOUSEHOLDER}(e,f)$
$g_{i}\leftarrow h_{b}h_{a}(1\oplus\text{ORTHOGONAL}(N-1,i_{r}))$
return
It is $\mathcal{O}(N^{3})$ because it has $N$ recursive steps, during each of which it takes $\mathcal{O}(N)$ to find the Householder reflection that maps $(1,0,\ldots,0)^{T}\in\mathbb{F}_{2}^{N}$ to the (odd-parity) binary expansion corresponding to the leading $N-1$ bits of $1\leq i\leq|\operatorname{O}(2n)|$. Applying these reflections to said basis vector (and the other $\mathcal{O}(N)$ ones from the previous recursion steps) then takes $\mathcal{O}(N)$ steps each, since doing so only involves computing inner products and adding vectors. Overall it therefore requires $\mathcal{O}(N^{2})$ time to apply the Householders to all columns. Uniqueness of this mapping follows from the ability to perform all these steps in reverse, i.e. by reading the binary representation of the corresponding index $i$ off the the leading column vectors at each inverse recursion step.